# Affine Chord Sobolev Inequalities from Integral Geometry

### v1

@@ 4. The star-shaped set ${\rm L}_{\alpha}f$
Building on the approach of Haddad and Ludwig [25, 26, 27], we introduce the following anisotropic version of $I_{\alpha+1}(f)$. Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. We consider
@@
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt.$
@@
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\rho_{K}(tu)^{n-\alpha}t^{n-1}\min\{f(y),f(y+tu)\}dtdudy$
$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(y),f(y+tu)\}dydtdu$
@@
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f),$ (4.1)
for measurable $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ and star-shaped set $K\subset\mathbb{R}^{n}$ with measurable radial function. In particular,
@@
$=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}t^{n-1}\min\{f(x),f(x+tu)\}dxdtdu$
@@
$\sup\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
@@
$\inf\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
@@
If $f:\mathbb{R}^{n}\to[0,\infty)$ is log-concave and in $L^{1}(\mathbb{R}^{n})$, then ${\rm L}_{\alpha}f$ is a convex set for every $\alpha>0$.
@@
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}t^{\alpha-1}g(tu)dt.$
Applying a result of Ball [5, Theorem 5], we obtain that
$u\mapsto\left(\int_{0}^{\infty}t^{\alpha-1}g(tu)dt\right)^{-\frac{1}{\alpha}}$
defines a norm on $\mathbb{R}^{n}$, which concludes the proof. ∎
5. Affine isoperimetric inequalities for $\alpha\in(0,n)$
In this section, we aim to establish an affine version of Theorem 1. To this end, we first prove the following crucial lemma, which uses the Riesz rearrangement inequality. We say that $A\subset\mathbb{R}^{n}$ is homothetic to $B\subset\mathbb{R}^{n}$ if there exist $x\in\mathbb{R}^{n}$ and $a>0$ such that $A=x+aB$.
@@ 5. Affine isoperimetric inequalities for $\alpha\in(0,n)$
The equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost all $r>0$, the level set $\{f\geq r\}$ has measure zero or is homothetic to $D$ up to a set of measure zero.
@@
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\{f\geq r\}}(y)}{\|x-y\|_{K}^{q}}dxdydr$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}(y)dxdydtdr.$
@@
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}$ $(y)dxdy$
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq r\}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq r\}}(y)dxdy,$
@@
If the equality holds, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}$ $(y)dxdy$
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy$
for almost all $(r,t)\in(0,\infty)^{2}$. For sufficiently large $t>0$, the assumptions of Theorem 2.2 are fulfilled. Therefore, up to sets of measure zero, we have
$\{f\geq r\}=x+aD,~~t^{-1/q}K=bD,$
where $D$ is an ellipsoid and $x\in\mathbb{R}^{n}$. Since $K=(\frac{|K|}{|D|})^{1/n}D$, the ellipsoid $D$ is not dependent on $t$ and $r$, which concludes the proof. ∎
@@
Let $\alpha\in(0,n)$. For non-negative $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$,
@@
If $|{\rm L}_{\alpha}f^{\star}|<\infty$, there is equality if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi$.
@@
$\lim_{k\rightarrow\infty}\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f_{k}(x),f_{k}(x+tu)\}dxdt=\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt,$
@@
$\lim_{k\rightarrow\infty}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{$ $f_{k}(x),f_{k}(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}du$
$=\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}t^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+tu)\}dxdt\Big)^{\frac{n}{\alpha}}du.$
@@
Since $f\in L^{n/(n+\alpha)}(\mathbb{R}^{n})$ and $f_{k}$ has compact support, $|{\rm L}_{\alpha}f_{k}|<\infty$. By the definition of the Schwarz symmetral, we have $(f_{k})^{\star}\leq f^{\star}$. Then the first part of the proof implies that
@@
For the second inequality, noting that $\|f\|_{\frac{n}{n+\alpha}}=\|f^{\star}\|_{\frac{n}{n+\alpha}}$ and ${\rm L}_{\alpha}f^{\star}$ is a ball, we obtain from Theorem Theorem 1, (4.1), and Lemma 5.2 that
@@
If equality holds throughout, it follows from Theorem Theorem 1 that $f=c\operatorname{1}_{B}$ with $B$ a ball. By Lemma 5.2, $f=f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. This characterizes the equality cases. ∎
6. Affine isoperimetric inequalities for $\alpha>n$
As mentioned in the introduction, Theorem 1 extends the classical isoperimetric inequality for chord power integrals to function. It also provides an $L^{1}$ analogue of the HLS inequalities, for which both inequalities reverse when $\alpha>n$. In our setting, however, this reversal does not appear in such a simple form here.
Before addressing this nontrivial case, we first present the following lemma, which serves as a counterpart to Lemma 5.1
@@ 6. Affine isoperimetric inequalities for $\alpha>n$
Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost $r>0$, the level set $\{f\geq r\}$ has measure zero or is homothetic to $D$ up to sets of measure zero.
@@
$\|z\|_{K}^{q}=\int_{0}^{\infty}\operatorname{1}_{\mathbb{R}^{n}\setminus t^{1/q}K}(z)dt.$
@@
$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\|x-y\|_{K}^{q}\operatorname{1}_{\{f\geq r\}}(y)dxdydr$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{\mathbb{R}^{n}\setminus t^{1/q}K}(x-y)\operatorname{1}_{\{f\geq r\}}(y)dxdydrdt$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)(1-\operatorname{1}_{t^{1/q}K}(x-y))\operatorname{1}_{\{f\geq r\}}(y)dxdydrdt.$
@@
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)\operatorname{1}_{t^{1/q}K}(x-y)\operatorname{1}$ ${}_{\{f\geq r\}}(y)dxdy$ (6.2)
$\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}^{\star}}(x)\operatorname{1}_{t^{1/q}K^{\star}}(x-y)\operatorname{1}_{\{f\geq r\}^{\star}}(y)dxdy.$
Note that $\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq r\}}(x)dx<\infty$ for almost all $r>0$, as $f\in L^{1}(\mathbb{R}^{n})$. Combing (6.2) with the calculation in (6.1), we have
@@
Moreover, if equality holds in (6.3), then there exists a null set $N\subset(0,\infty)^{2}$ such that equality holds in (6.2) for all $(r,t)\in(0,\infty)^{2}\setminus N$.
For almost every $t\in(0,\infty)$, we have $(r,t)\in(0,\infty)^{2}\setminus N$ for almost every $r>0$. For $t$ sufficiently small, Theorem 2.2 guarantees the existence of a centered ellipsoid $D$ and $x\in\mathbb{R}^{n}$ such that
$\{f\geq r\}=x+aD\quad\text{and }\quad t^{1/q}K=bD.$
@@
With the previous lemma established, for $\alpha>n$, we have
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy$ $\geq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f^{\star}(x),f^{\star}(y)\}}{|x-y|^{n-\alpha}}dxdy$ (6.4)
$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f^{\star}\geq r\}}(x)\operatorname{1}_{\{f^{\star}\geq r\}}(y)}{|x-y|^{n-\alpha}}dxdy$
$=\sigma_{n,\alpha}\int_{0}^{\infty}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
However, Lemma 3.2 is not applicable in this case because the inequality holds in the opposite direction. We therefore apply a different approach to obtain the result as follows.
Theorem 6.2.
Let $\alpha>n$ and $f$ be a non-zero, non-negative function in $L^{1}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$. Then
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$
Equality holds if and only if $f=c\operatorname{1}_{B}$ with $B$ a ball and $c>0$.
Proof.
As we calculated in (6.4),
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy\geq\sigma_{n,\alpha}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr.$
By Jensen’s inequality, we have
$\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|^{\frac{n+\alpha}{n}}dr\geq\Big(\frac{1}{\|f\|_{\infty}}\int_{0}^{\|f\|_{\infty}}|\{f\geq r\}|dr\Big)^{\frac{n+\alpha}{n}}=\|f\|_{\infty}^{-\frac{n+\alpha}{n}}\|f\|_{1}^{\frac{n+\alpha}{n}},$ (6.5)
and thus
$\sigma_{n,\alpha}\|f\|_{1}^{\frac{n+\alpha}{n}}\|f\|_{\infty}^{-\frac{\alpha}{n}}\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy.$ (6.6)
If equality holds in (6.6), then equality must occur in the application of Jensen’s inequality in (6.5). This in turn implies that $|\{f\geq r\}|$ is constant for almost every $r\in(0,\|f\|_{\infty})$. Equivalently, this happens if and only if $f=c\operatorname{1}_{E}$, where $E$ is a measurable set of finite measure and $c>0$.
When $f=c\operatorname{1}_{E}$, the inequality (6.6) reduces to the geometric case:
$I_{\alpha+1}(E)\geq\Big(\frac{I_{\alpha+1}(B^{n})}{\omega_{n}^{(n+\alpha)/n}}\Big)|E|^{\frac{n+\alpha}{n}},$
with equality if and only if $E$ is a ball. This then completes the characterization of the equality cases. ∎
@@
By letting $q=n-\alpha$ in Lemma 6.2 and combining it with (4.1), we obtain
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{|x-y|^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\geq n\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
@@
$|{\rm L}_{\alpha}f|$ $\geq\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
@@
Building on the results established above, we proceed to the proof of Theorem 3.
@@
For the first inequality, if the left side is $\infty$, the inequality holds trivially. We then assume that
@@
For the second inequality, we may assume that $|{\rm L}_{\alpha}f|$ is finite. Note that ${\rm L}_{\alpha}f^{\star}$ is a ball. Then Lemma 6.3 and (4.1) implies that
@@
Combining Theorem 6.2 for the right-hand side with the facts $\|f\|_{1}=\|f^{\star}\|_{1}$ and $\|f\|_{\infty}=\|f^{\star}\|_{\infty}$, we deduce
@@
The equality characterization is given by Theorem 6.2 and Lemma 6.3. ∎
7. Radial mean bodies

### v2

@@ 3. The star-shaped set ${\rm L}_{\alpha}f$
Building on the approach of Haddad and Ludwig [24, 25, 26], we introduce the following anisotropic integral. Let $f:\mathbb{R}^{n}\rightarrow[0,\infty)$ be a measurable function. We consider
@@
$\rho_{{\rm L}_{\alpha}f}(u)^{\alpha}=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr.$
@@
$=\int_{\mathbb{R}^{n}}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\rho_{K}(ru)^{n-\alpha}r^{n-1}\min\{f(y),f(y+ru)\}drdudy$
$=\int_{\mathbb{S}^{n-1}}\rho_{K}(u)^{n-\alpha}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(y),f(y+ru)\}dydrdu$
@@
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f).$ (3.1)
In particular,
@@
$=\frac{1}{n}\int_{\mathbb{S}^{n-1}}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}r^{n-1}\min\{f(x),f(x+ru)\}dxdrdu$
@@
$\sup\Big\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\Big\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
@@
$\inf\Big\{\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy:K\subset\mathbb{R}^{n}~\text{star-shaped},|K|=\omega_{n}\Big\}=n\omega_{n}^{\frac{n-\alpha}{n}}|{\rm L}_{\alpha}f|^{\frac{\alpha}{n}}.$
@@
Let $\alpha>0$. If $f:\mathbb{R}^{n}\to[0,\infty)$ is a log-concave, integrable function with compact support, then ${\rm L}_{\alpha}f$ is a convex body.
@@
$\rho_{{\rm L}_{\alpha}f}(u)=\Big(\int_{0}^{\infty}r^{\alpha-1}g(ru)dr\Big)^{1/\alpha}.$
It follows from [18, Corollary 4.2] that ${\rm L}_{\alpha}f$} is a convex body. ∎
4. Affine chord Sobolev inequalities
In this section, we establish the affine chord Sobolev inequalities stated in Theorem 1 and Theorem 2. To this end, we first prove a crucial lemma based on the Riesz rearrangement inequality.
Recall that a set $A\subset\mathbb{R}^{n}$ is homothetic to $B\subset\mathbb{R}^{n}$ if there exist $x\in\mathbb{R}^{n}$ and $a>0$ such that $A=x+aB$. Moreover, $A$ is said to be equivalent to $B$ if their symmetric difference $A\triangle B$ is a null set. The following lemma provides a general form of the Riesz rearrangement inequality; see also [10, 24, 25, 26].
@@ 4. Affine chord Sobolev inequalities
Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost all $t>0$, the level set $\{f\geq t\}$ has measure zero or is homothetic to $D$ up to a set of measure zero.
@@
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{q}}dxdy$ $=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{\{f\geq t\}}(y)}{\|x-y\|_{K}^{q}}dxdydt$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{r^{-1/q}K}(x-y)\operatorname{1}_{\{f\geq t\}}(y)dxdydrdt.$
@@
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{r^{-1/q}K}$ $(x-y)\operatorname{1}_{\{f\geq t\}}(y)dxdy$ (4.1)
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq t\}}(x)\operatorname{1}_{r^{-1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq t\}}(y)dxdy,$
@@
If equality holds, then (4.1) is an equality for almost all $(t,r)\in(0,\infty)^{2}$. For sufficiently large $r>0$, the assumptions of Theorem 2.2 are satisfied. Consequently, up to sets of measure zero, we have
$\{f\geq t\}=x+aD,\qquad r^{-1/q}K=bD,$
where $a,b>0$, $x\in\mathbb{R}^{n}$, and $D$ is a centered ellipsoid with the same volume as $K$. It follows that $K=\big(\frac{|K|}{|D|}\big)^{1/n}D$, showing that $D$ is independent of $t$ and $r$, which completes the proof. ∎
The next lemma shows that, for $\alpha\in(0,n)$, the volume of ${\rm L}_{\alpha}f$ does not increase under Schwarz symmetrals.
@@
Let $\alpha\in(0,n)$. For non-negative $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$,
@@
If $|{\rm L}_{\alpha}f^{\star}|<\infty$, there is equality if and only if $f$ is a translate of $f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$.
@@
$\lim_{k\rightarrow\infty}\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f_{k}(x),f_{k}(x+ru)\}dxdr=\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr,$
@@
$\lim_{k\rightarrow\infty}\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{$ $f_{k}(x),f_{k}(x+ru)\}dxdr\Big)^{\frac{n}{\alpha}}du$
$=\int_{\mathbb{S}^{n-1}}\Big(\int_{0}^{\infty}r^{\alpha-1}\int_{\mathbb{R}^{n}}\min\{f(x),f(x+ru)\}dxdr\Big)^{\frac{n}{\alpha}}du.$
@@
Since $f\in L^{\frac{n}{n+\alpha}}(\mathbb{R}^{n})$ and $f_{k}$ has compact support, $|{\rm L}_{\alpha}f_{k}|<\infty$. By the definition of the Schwarz symmetral, we have $(f_{k})^{\star}\leq f^{\star}$. Then the first part of the proof implies that
@@
For the second inequality, noting that $\|f\|_{\frac{n}{n+\alpha}}=\|f^{\star}\|_{\frac{n}{n+\alpha}}$ and that ${\rm L}_{\alpha}f^{\star}$ is a ball, we obtain from the chord Sobolev inequality (1.3), (3.1), and Lemma 4.2 that
@@
If equality holds throughout, it follows from the chord Sobolev inequality that $f^{\star}=c\operatorname{1}_{B}$, where $B$ is a ball. By Lemma 4.2, we have $f=f^{\star}\circ\phi^{-1}$ for some volume-preserving linear transformation $\phi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. This characterizes the equality cases. ∎
For $\alpha>n$, we state the following analogue of Lemma 4.1, which also generalizes [4, Lemma 3.2]. The proof follows a similar approach and is included for completeness.
@@
Equality holds if and only if $K$ is equivalent to a centered ellipsoid $D$ and for almost $t>0$, the level set $\{f\geq t\}$ has measure zero or is homothetic to $D$ up to sets of measure zero.
@@
$\|z\|_{K}^{q}=\int_{0}^{\infty}\operatorname{1}_{\mathbb{R}^{n}\setminus r^{1/q}K}(z)dr.$
@@
$=\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\|x-y\|_{K}^{q}\operatorname{1}_{\{f\geq t\}}(y)dxdydt$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{\mathbb{R}^{n}\setminus r^{1/q}K}(x-y)\operatorname{1}_{\{f\geq t\}}(y)dxdydtdr$
$=\int_{0}^{\infty}\int_{0}^{\infty}\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)(1-\operatorname{1}_{r^{1/q}K}(x-y))\operatorname{1}_{\{f\geq t\}}(y)dxdydtdr.$
@@
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)\operatorname{1}_{r^{1/q}K}(x-y)$ $\operatorname{1}_{\{f\geq t\}}(y)dxdy$ (4.4)
$\leq\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f^{\star}\geq t\}}(x)\operatorname{1}_{r^{1/q}K^{\star}}(x-y)\operatorname{1}_{\{f^{\star}\geq t\}}(y)dxdy.$
Note that $\int_{\mathbb{R}^{n}}\operatorname{1}_{\{f\geq t\}}(x)dx<\infty$ for almost all $t>0$, as $f\in L^{1}(\mathbb{R}^{n})$. Combing (4.4) with the calculation in (4.3), we have
@@
Moreover, if equality holds in (4.5), then there exists a null set $N\subset(0,\infty)^{2}$ such that equality holds in (4.4) for all $(t,r)\in(0,\infty)^{2}\setminus N$. For almost all $r\in(0,\infty)$, we have $(t,r)\in(0,\infty)^{2}\setminus N$ for almost all $t>0$. For $r$ sufficiently small, Theorem 2.2 guarantees the existence of a centered ellipsoid $D$ and $x\in\mathbb{R}^{n}$ such that
$\{f\geq t\}=x+aD\quad\text{and }\quad r^{1/q}K=bD.$
@@
Similarly, for $\alpha>n$, we show that $|{\rm L}_{\alpha}f|$ does not decrease under Schwarz symmetrals.
@@
By letting $q=\alpha-n$ in Lemma 4.3 and combining it with (3.1), we obtain
$\int_{\mathbb{R}^{n}}\int_{\mathbb{R}^{n}}\frac{\min\{f(x),f(y)\}}{\|x-y\|_{K}^{n-\alpha}}dxdy=n\tilde{V}_{\alpha}(K,{\rm L}_{\alpha}f)\geq n\tilde{V}_{\alpha}(K^{\star},{\rm L}_{\alpha}f^{\star}).$
@@
$|{\rm L}_{\alpha}f|$ $=\tilde{V}_{\alpha}({\rm L}_{\alpha}f,{\rm L}_{\alpha}f)$
@@
Building on the preceding lemmas, we now establish the affine chord Sobolev inequalities for $\alpha>n$.
@@
If the right-most side is $\infty$, the inequality holds trivially. We then assume that
@@
For the first inequality, we may assume that $|{\rm L}_{\alpha}f|$ is finite. Note that ${\rm L}_{\alpha}f^{\star}$ is a ball. Then Lemma 4.4 and (3.1) imply that
@@
Combining the chord Sobolev inequality (1.4) with $\|f\|_{1}=\|f^{\star}\|_{1}$ and $\|f\|_{\infty}=\|f^{\star}\|_{\infty}$, we deduce
@@
The equality case follows from the optimizers of (1.4) together with Lemma 4.4. ∎
5. Radial mean bodies for log-concave functions